OpenStudy (anonymous):

Two planes leave from Topeka, Kansas. The first plane travels east at a rate of 278 mph. The second plane travels west at a speed of 310 mph. How long will it take for them to be 1176 miles apart? 2 hours 4 hours 6 hours 8 hours

OpenStudy (anonymous):

in 1 hour both travel (278+310)m=588m to travel 558m, they need 1 hr to travel 1176m, they need 1176/588 hrs = 2 hrs

OpenStudy (anonymous):

r t = d 278 t + 310 t = 1176 t = 2 hours